OpenStudy (deadpoolroxz):

Which answer shows this equation in standard form? 8 – 4(x – y) = –6x + 6 A. –10x + 4y = –2 B. 2x + 4y = –2 C. 2x – 4y = –2 D. 2x + 4y = 2

OpenStudy (deadpoolroxz):

@jordanloveangel

OpenStudy (deadpoolroxz):

@MAli13chineta

OpenStudy (deadpoolroxz):

supz jordan

OpenStudy (jordanloveangel):

ayeee!!

OpenStudy (deadpoolroxz):

thanxz jessie wait like demmi lavato ?

OpenStudy (jordanloveangel):

8 - 4(x - y) = 2x + 6 8 - 4x +4y = 2x + 6 2 = 6x -4y so ur answer is 6x - 4y = 2 but Got -6x + 4y = -2. That option isn't available so what would be the right answer???

OpenStudy (deadpoolroxz):

yeah thanx bye

OpenStudy (jordanloveangel):

hi bye :(

OpenStudy (deadpoolroxz):

bye